If f(x) = 2x^2 - 7x + 3 find: f(2) and f(-3)
Answer by oscargut(2103)

If f(x) = 2x^2 - 7x + 3 find: f(2) and f(-3)
f(2)= 2(2)^2-7(2)+3=8-14+3=-3
f(-3)= 2(-3)^2-7(-3)+3=18+21+3=42